OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (anonymous):

Expand right side using your distributive property, gather like terms and isolate for p.

jhonyy9 (jhonyy9):

-6p+7=6p-9+40-16p 7+9-40=6p+6p-16p -24=-4p p=-24/(-4) = 6 p=6

OpenStudy (anonymous):

thank you so much, I was out of school for like 6 months due to my cancer treatment so Trig makes NO sense at the moment